Zero Turbulence Airline provides air transportation services between Los Angeles, California; and Kona, Hawaii. A single Los Angeles to Kona round-trip flight has the following operating statistics:

Fuel $13,395

Flight crew salaries 10,260

Airplane depreciation 4,845

Variable cost per passenger—business class 60 Variable cost per passenger—economy class 50 Round-trip ticket price—business class 570 Round-trip ticket price—economy class 310 It is assumed that the fuel, crew salaries, and airplane depreciation are fixed, regardless of the number of seats sold for the round-trip flight. If required round the answers to nearest whole number.

a. Compute the break-even number of seats sold on a single round-trip flight for the overall product, E. Assume that the overall product is 10% business class and 90% economy class seats. Total number of seats at break-even ___ seats

b. How many business class and economy class seats would be sold at the break-even point? Business class seats at break-even ___ seats Economy class seats at break-even ___ seats
Answer #1

Fixed cost = 13395+10260+4845 = 28500

Weighted average contribution margin per unit = (570-60*10%)+(310-50*90%) = 285

a) Break even point = Fixed cost/Weighted average contribution margin per unit = 28500/285 = 100 seats

b) Business class seats at break even = 100*10% = 10 seats

Economy class seats at break even = 100*90% = 90 seats